OpenStudy (anonymous):

A rectangle has a length that is 3 feet longer than its width. its perimeter is 26 feet. What is the width of the rectangle?

OpenStudy (solomonzelman):

w= x+3 l= x (you would need to add 3 to the width to make it equal to the length, THAT'S WHY w=x+3 and l=x ) P=2(w+l) in your case, 26 =2( x + [x+3] ) Solve for x, and then tell me the width (that's equal to x+3, but tell me the width not in terms of x, rather as value)

OpenStudy (anonymous):

10?

OpenStudy (solomonzelman):

26 =2( x + [x+3] ) 26 =2( 2x+3 ) 26 =4x+6 20=4x x=5 w =x+3=5+3 =8 l =x = 5 Checking: 26=2(8+5) 26=2(13) 26=26

OpenStudy (anonymous):

Thanks

OpenStudy (solomonzelman):

You welcome !